# On the inadequacy of nudging data assimilation algorithms for non-dissipative systems

### v2

@@ 3.1. Preliminaries and Analytical Results
We note that here $H^{0}_{\text{per}}=L^{2}_{\text{per}}$, with $L^{2}_{\text{per}}$ denoting the space of $L^{2}$ functions subject to $L$-periodic boundary conditions.
@@
$v_{t}=F(v)+\mu I_{h}(u-v).$ (18)
Here $I_{h}$ is a linear spatial interpolant with associated length scale $h$ and $\mu$ is a constant feedback-control nudging parameter. For our purposes we will consider only $I_{h}\mathrel{\mathop{\ordinarycolon}}=P_{M}$, where $P_{M}$ is given to be a projection operator onto the lowest $M$ Fourier modes, i.e. $h\sim\frac{L}{M}$, where $L$ is the period for the underlying evolutionary equation Equation˜14. As we require $P_{M}u$ to be well defined we will only consider solutions to Equation˜14 that are in $L^{2}_{\text{per}}$ for all positive time.
For the purpose of this study we will let $F$ be given such that we obtain the KdV equation:
@@
Here $u\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\times[0,\infty)\to\mathbb{R}$ represents the amplitude of a wave and $\delta$ is the dispersive coefficient. We consider the KdV equation equipped with the periodic boundary condition $u(x,t)=u(x+2,t)$, for all $x\in\mathbb{R}$ and all time $t\geq 0$ with spatial mean zero. Hence, the period in this case is $L=2$. We note that choosing the length of the spatial domain to be 2 is arbitrary and was chosen as this choice of domain appears in the literature, see, e.g., [80].
@@
We note that below and through the remainder of the paper, we will use the notation $\hat{f}$ to reference the Fourier modes of an arbitrary function $f$ in $L^{2}_{\text{per}}$. That is,
@@
Let $k\in\mathbb{Z}\setminus\{0\}$ and $\phi\in L^{2}_{\text{per}}$ with $\phi$ having spatial mean zero. The following hold true:
@@
Suppose $k\in\mathbb{Z}\setminus\{0\}$ and $u^{in},f\in H^{2}_{\text{per}}$ such that $f(x+\frac{L}{k})=f(x)$ and $u^{in}(x+\frac{L}{k})=u^{in}(x)$ for all $x\in\mathbb{R}$ and $u^{in}$ and $f$ both have spatial mean zero. Let $u(x,t)$ be the corresponding solution of the damped and driven KdV equation:
@@
Suppose $k\in\mathbb{Z}\setminus\{0\}$ is given and suppose $u^{in}_{j}\in H^{2}_{\text{per}}$, for $j=1,2$, satisfying
@@
We note that as Equation˜24 is a particular case of Equation˜24 (i.e. $\gamma=0,\beta=0$), then (i) follows immediately from Proposition˜3.3. Part (ii) then follows immediately from part (i) combined with Proposition˜3.2.
Let us now prove part (iii). Let $u_{1}$ and $u_{2}$ be two solutions to Equation˜19 under the assumptions of the proposition. Without loss of generality, suppose $\mathinner{\!\left\lVert u_{1}^{\text{in}}\right\rVert}_{L^{2}}>\mathinner{\!\left\lVert u_{2}^{\text{in}}\right\rVert}_{L^{2}}$. We note that the $L^{2}$ norm is a conserved quantity for Equation˜19 and thus for $j=1,2$ we have that $\mathinner{\!\left\lVert u_{j}(t)\right\rVert}_{L^{2}}=\mathinner{\!\left\lVert u_{j}^{\text{in}}\right\rVert}_{L^{2}}$ holds for all $t\in\mathbb{R}$. The result follows from an application of the reverse triangle inequality as
@@
Proposition˜3.5 can be adjusted without change to allow $u_{1}$ and $u_{2}$ to be initialized with periods $\frac{L}{k_{1}}$ and $\frac{L}{k_{2}}$, respectively. The only changes in this case are that $u_{j}$ remain $\frac{2}{k_{j}}$ periodic for all $t\in\mathbb{R}$ in (i) and $M<\min\{\mathinner{\!\left\lvert k_{1}\right\rvert},\mathinner{\!\left\lvert k_{2}\right\rvert}\}$ is required for (ii).
@@
The proof of this follows straightforwardly from Proposition˜3.5. Assume, by way of contradiction, that the KdV equation enjoys the finitely many determining Fourier modes property in $L^{2}_{\text{per}}$. That is, there exists some $M\in\mathbb{N}$ such that for any two solutions $u_{1}$ and $u_{2}$ to Equation˜19 that if
@@
Now, we will simply use $k=M+1$, and consider $u_{1}$ and $u_{2}$ to be any two solutions to Equation˜19 such that $u_{1}$ and $u_{2}$ both are smooth enough and have zero spatial mean, are $\frac{L}{k}$ periodic initially, and $\mathinner{\!\left\lVert u_{1}^{in}\right\rVert}_{L^{2}}\neq\mathinner{\!\left\lVert u_{2}^{in}\right\rVert}_{L^{2}}$. We note that two such solutions must exist, as one could take $u_{1}$ to be the zero solution and $u_{2}$ to be the solution generated from the initial data $u_{2}^{in}=\cos(\frac{2\pi}{L}kx).$
@@
The proof above can be adapted almost without change to the unforced 2D incompressible Euler equations and other non-dissipative systems with a similar nonlinearity. This is due to the fact that the convolutional structure of the Fourier representation of the nonlinear term is the same in both cases and that the Euler equations conserve the solenoidal component of the $L^{2}$ norm for smooth initial data.
In a future work, we will extend this argument to the damped and driven Euler equations, which introduce additional complexities due to the external forcing and damping terms.
@@ 3.2. Computational Results
In this section we will detail the numerical methods we utilize to simulate Equations˜19 and 20. All simulations depicted in this study were performed Matlab (R2023a) using our own code. For this computational study we implemented the KdV equation on $\mathbb{R}$, subject to periodic B.C. with period 2 with pseudo-spectral methods. Specifically, we implement the equations using an integrating factor method to compute the linear term exactly with a fourth-order Runge-Kutta scheme for the timestepping with the nonlinear term computed explicitly. For an overview of integrating factor schemes see e.g. [55, 78] and the references contained within. Our particular choice of numerical scheme, which is given Algorithm˜1, was chosen as it allows for larger timesteps than traditional schemes for KdV, such as the Zabusky-Kruskal explicit leapfrog scheme seen in the literature [80].
We note that in the simulations of Equation˜19, depicted below, we utilized different spatial resolutions depending on the particular choice of initial profile and dispersion coefficient, $\delta^{2}$. We ensured in our simulations that $\Delta t\lesssim{\Delta x}^{2}$, which was consistent with the implementation of this timestepping scheme in chapter 10 of [78] and appeared to be numerically stable for all spatial resolutions we utilized in this work. Here $\Delta t=t_{n+1}-t_{n}$ and $\Delta x=\frac{2}{N}$ represent the temporal and spatial resolution scales with $N$ indicating the spatial resolution, given in powers of $2$. While we do not explicitly state the spatial resolution for each simulation, one can see in the spectral plots, i.e. Figures˜4, 5, 6, 7, 8 and 9, that the x-axis covers $[1,N/2]$. In this study, $N$ ranges between $2^{7}$ and $2^{11}$, depending on the amount of Fourier modes required to accurately simulate a given solution. $N$ is chosen specifically to ensure that the solution is well resolved, which here means that Fourier coefficients of the reference solution decay to machine precision before the dealiasing cutoff, i.e. the vertical red line featured in Figures˜4, 5, 6, 7, 8 and 9. That is, the momentum decays exponentially in wave number, we take $N$ large enough such that the momentum decays to machine precision (given by 2.216e162.216e-16) before the wave number $\frac{N}{3}$. We utilize 2/3 truncation for dealiasing to accurately calculate the nonlinear term, i.e., we set the last 1/3 of the Fourier modes to zero to prevent aliasing effects. Additionally, we note that while the red vertical line on spectral plots (see Figures˜9(b), 9(d), 7(b) and 8(b) denotes the dealiasing cutoff, the vertical blue line denotes the observational window cutoff, that is, the Fourier modes to the left of and including the blue vertical line are accurately observed and used in the computation of the nudging term.
@@
The nudging term is implemented numerically in Algorithm˜1 as an explicit term. This implementation of the feedback-control term that we used induces a CFL constraint required for the numerical scheme to remain stable, $\mu\lesssim\frac{2}{\Delta t}$. This CFL constraint can be removed by utilizing an implicit formulation of the nudging term, which we do not consider in this work. It is worth noting that the nudging term cannot be implemented numerically using multi-stage methods, such as RK4, due to the need to interpolate in time (see e.g. [67]). The simulations featured in this work primarily use a value of $\mu=100$, which was found to be sufficient for this study. For an in-depth examination of the role the value of the nudging parameter plays in convergence, see [15].
To test the effectiveness of nudging on the KdV equation Equation˜19 we utilized an “identical twin” experimental design. This experimental design is standard in the validation of data assimilation methods, where we first initialize a reference solution. This solution will be treated as the true solution which we will attempt to recover using observations. The observations will be given by the lowest $M$ Fourier modes. We note that the observations are considered to be perfect with no distortions (i.e., no error measurements) or stochastic noise.
We note that typically in the data assimilation literature for dissipative evolutionary equations, the equations are evolved forward in time long enough such that the solution remains in an absorbing set in a relevant phase space. That is, we run solutions long enough forward in time such that they are uniformly bounded with respect to time in $L^{2}$, $H^{1}$, or another relevant norm for all future times, see e.g., [74, 76] for a full definition of absorbing sets as well as their properties. In this study we do not evolve the equations forward in time to arbitrarily large times, as solutions are not dissipative and so there is not necessarily an absorbing set that attracts all bounded solutions after a long enough time. In fact, as the $L^{2}$ norm is a conserved quantity for Equation˜19, such an absorbing set cannot exist in $L^{2}$. We opted to simply evolve forward the initial profile a given amount of time, typically $t=1$, in order to ensure that the reference solution has developed interesting dynamics outside the observation domain. While solutions can develop small length scales not captured by the observed modes, they typically have a finite number of active Fourier modes, with the total number of modes active at a given time dependent on the initial condition, the current time, the total momentum in the system, and the dispersion parameter, $\delta^{2}$. An active Fourier mode in this sense means that $\mathinner{\!\left\lvert\hat{u}_{k}\right\rvert}>\varepsilon$, where $k\in\mathbb{Z}$ is the wavemode index and $\varepsilon$ indicates machine precision.
@@ 3.2.1. Classical KdV equation
In this section we will examine how the data assimilation nudging algorithm fails for Equation˜19. Nudging appears to fail subtly, with the extent to which it works appears to be determined by whether the active Fourier modes of the initial profile are contained in the annulus of the observed Fourier modes, as we will see later on in Figure˜8. We will start by considering examples such as in Proposition˜3.5 to illustrate how nudging fails.
@@
Using initial data of the form Equation˜29 we can now construct solutions to Equation˜19 which the nudging algorithm Equation˜20 explicitly fails to recover. Given a number of observed modes $M$, we can set $k_{0}\in\mathbb{Z}$ such that $\mathinner{\!\left\lvert k_{0}\right\rvert}\geq M+1$. Again, we emphasize that solutions initialized in this way (see Equation˜29) are initially $\frac{2}{k_{0}}$ periodic and remain $\frac{2}{k_{0}}$ periodic due to Proposition˜3.3. Noting that $M<k_{0}$, it follows immediately from Proposition˜3.5 that $P_{M}u(t)\equiv 0$ where $u$ is the solution of Equation˜19 generated from the initial data prescribed by Equation˜29. This shows that a fixed number of observed Fourier modes are insufficient to distinguish the zero solution from solutions initialized with suitable high frequency oscillations, and so one should not expect data assimilation to succeed. Indeed, if we choose $v$ to be the solution of Equation˜20 initialized by the identically zero solution, then it is easy to deduce that $P_{M}(u-v)\equiv 0$ holds for all time and so the solution to Equation˜20 simply corresponds to a solution to Equation˜19 albeit with a different initial condition than the reference solution.
@@
One can see computationally in Figure˜4 that nudging fails to recover the solution $u(x,t)$, of Equation˜19 generated from the initial data prescribed by Equation˜29. In Figure˜4 the solution, $v(x,t)$, of Equation˜20 is initialized with the identically zero solution. Exactly as one expects, we find that $v\equiv 0$ holds for the entire simulation. Similarly, we see in Figure˜5 the roles of $u$ and $v$ are reversed. That is, we instead initialize the solution to $v$ with Equation˜29 and the reference solution, $u$, is initialized with the identically zero solution. We see in Figure˜5 similar behavior occurs, as the feedback-control term remains zero for the entire simulation and both equations evolve as solutions of Equation˜19 with different initial conditions. While this is expected, it is nevertheless an interesting result as it shows that the initialization of the assimilated solution to Equation˜20 matters. This fact is fundamentally different to many analytical results regarding the application of the nudging algorithm to dissipative equations, as for dissipative systems the effect of the initial condition fades given enough time, contrary to what occurs in this case.
@@
The previous counterexample illustrates that data assimilation methods cannot uniquely determine the observed reference solution when the observations are identically zero, but one may hope that more physically realistic (i.e. not identically zero) observations will yield better results. This is true to some extent, however the level of success is highly dependent on the initial profile of the reference solution and one can similarly construct solutions that exhibit high frequency oscillations that nudging cannot recover. If given an analytic solution $u$ in $L^{2}_{\text{per}}((0,L))$ of Equation˜19, we can indeed construct infinitely many solutions where nudging fails to synchronize with the reference solution given a finite number of observed Fourier modes.
@@
By Parseval’s identity $\mathinner{\!\left\lVert u\right\rVert}_{L^{2}}^{2}=2\sum_{k\in\mathbb{Z}}\mathinner{\!\left\lvert\hat{u}_{k}\right\rvert}^{2}=K^{2}$. Here $\varepsilon$ is a specified constant, for computations we will use $\varepsilon\sim e-16$ to denote machine precision. The purpose of finding such an $M$ is to separate the wavemodes into active low modes and inactive high modes. At any fixed time $t>0$ such an $M$ is guaranteed to exist as $u$ is analytic and therefore the Fourier coefficients must decay exponentially fast in wavenumber. In our computations we run only for a finite amount of time, so we take $M$ for $u$ to be the supremum over all such $M$’s taken across all simulated times. Without loss of generality we will assume that observation window consists of all $M$ modes, and so essentially the entire solution $u$ is known down to machine precision. We note here that $M$ is vital to the construction we outline below, thus examples we construct work only for finite times. This is due to the fact that we require a uniform bound on the number of active wavemodes to extend this arbitrarily far in time.
Now, we can take the given solution $u$ and use it to generate artificial solutions for which nudging appears to fail. Simply consider the solution generated by the following initial data:
@@
Here $c_{1}$ and $c_{2}$ are positive constants chosen to enforce that $\mathinner{\!\left\lVert\tilde{u}^{in}\right\rVert}_{L^{2}}=K$ and $c_{2}>\varepsilon$. That is, we are modifying the initial profile to take some of the momentum out of the active modes and shifting it into to wavemode $2M+k$, where $k$ is an arbitrary positive integer. As $k$ is arbitrary, we can construct infinitely many such solutions, all of which have the same momentum as the given solution, $u$, but with a new active frequency in the initial profile. Note that the choice of using ${u}^{in}$ is arbitrary, and one can do this exact construction using the solution of $u$ evaluated at any given time.
We performed computational tests using the reference solution, $\tilde{u}$, of Equation˜19 with initial conditions prescribed by Equation˜31. Here we took a solution, $u$, of Equation˜19 which was initialized simply with $u^{in}=\cos(\pi x)$, which we found $M=50$ was suitable for the construction in Equation˜31. Once $M=50$ was determined, we simply utilized $\tilde{u}^{in}=\cos(\pi\cdot x)+\cos(100\cdot\pi\cdot x)$, which was renormalized to ensure that $\mathinner{\!\left\lVert\tilde{u}^{in}\right\rVert}_{L^{2}}=\mathinner{\!\left\lVert u^{in}\right\rVert}_{L^{2}}$. We see in Figure˜6 the result from applying nudging to attempt to recover the constructed reference solution $\tilde{u}$. Here the nudging solution, $v$, of Equation˜20 is initialized to be identically zero. We see in Figure˜6 that the low modes of $v$ synchronize with $\tilde{u}$ to a fixed level of precision, but the reference solution contains high frequency oscillations that $v$ never develops. The error in both the high and low modes appears approximately constant after a very brief initial decay in the observed modes. We note that in additional tests (not depicted here) the same behavior in the $L^{2}$ error was observed up to time $t=1000$ without any significant deviations in the error. One can clearly see activity in the frequencies immediately surrounding $k=100$ and $k=200$ that is simply not present in the nudged solution. Not only is this high frequency activity never generated, but nudging features no mechanism to generate targeted high mode oscillations. That is, if one knew a priori that the solution featured oscillations at wavemode $k=100$, but one still only had observations of the first $50$ wavemodes, nudging does not have any mechanism to encourage oscillations at any unobserved frequencies beyond the nonlinear term.

### v3

@@ 3.1. Preliminaries and Analytical Results
Here $H^{0}_{\text{per}}=L^{2}_{\text{per}}$, with $L^{2}_{\text{per}}$ denoting the space of $L^{2}$ functions subject to $L$-periodic boundary conditions. For simplicity of notation we will drop the subscript “per” as we will only be considering periodic functions in this work.
$v_{t}=F(v)+\mu I_{h}(u-v).$ (16)
Here $I_{h}$ is a linear spatial interpolant with associated length scale $h$ and $\mu$ is a constant feedback-control nudging parameter. For our purposes we will consider only $I_{h}\mathrel{\mathop{\ordinarycolon}}=P_{M}$, where $P_{M}$ is given to be a projection operator onto the lowest $M$ Fourier modes, i.e. $h\sim\frac{L}{M}$, where $L$ is the period for the underlying evolutionary equation Equation˜14. As we require $P_{M}u$ to be well defined we will only consider solutions to Equation˜14 that are in $L^{2}$ for all positive time.
An important property that is often required for the success of data assimilation algorithms is that the underlying system has finitely many determining modes. This property was first introduced in [37] for the 2D incompressible Navier-Stokes equations and later extended to other dissipative systems, see e.g. [54, 23, 36]. We now give the definition of this property.
@@
For the purpose of this study we will let $F$ result in the KdV equation:
@@
Here $u\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\times[0,\infty)\to\mathbb{R}$ represents the amplitude of a wave and $\delta$ is the dispersive coefficient. We consider the KdV equation equipped with the periodic boundary condition $u(x,t)=u(x+2,t)$, for all $x\in\mathbb{R}$ and all time $t\geq 0$ with spatial mean zero. Hence, the period in this case is $L=2$. Choosing the length of the spatial domain to be 2 is arbitrary and was chosen as this choice of domain appears in the literature, see, e.g., [80].
@@
Below and through the remainder of the paper, we will use the notation $\hat{f}$ to reference the Fourier modes of an arbitrary function $f$ in $L^{2}$. That is,
@@
Let $k\in\mathbb{Z}\setminus\{0\}$ and $\phi\in L^{2}$ with $\phi$ having spatial mean zero. The following hold true:
@@
Suppose $k\in\mathbb{Z}\setminus\{0\}$ and $u^{in},f\in H^{2}$ such that $f(x+\frac{L}{k})=f(x)$ and $u^{in}(x+\frac{L}{k})=u^{in}(x)$ for all $x\in\mathbb{R}$ and $u^{in}$ and $f$ both have spatial mean zero. Let $u(x,t)$ be the corresponding solution of the damped and driven KdV equation:
@@
Suppose $k\in\mathbb{Z}\setminus\{0\}$ is given and suppose $u^{in}_{j}\in H^{2}$, for $j=1,2$, satisfying
@@
As Equation˜24 is a particular case of Equation˜24 (i.e. $\gamma=0,\beta=0$), then (i) follows immediately from Proposition˜3.3. Part (ii) then follows immediately from part (i) combined with Proposition˜3.2.
Let us now prove part (iii). Let $u_{1}$ and $u_{2}$ be two solutions to Equation˜19 under the assumptions of the proposition. Without loss of generality, suppose $\mathinner{\!\left\lVert u_{1}^{\text{in}}\right\rVert}_{L^{2}}>\mathinner{\!\left\lVert u_{2}^{\text{in}}\right\rVert}_{L^{2}}$. We recall that the $L^{2}$ norm is a conserved quantity for Equation˜19 and thus for $j=1,2$ we have that $\mathinner{\!\left\lVert u_{j}(t)\right\rVert}_{L^{2}}=\mathinner{\!\left\lVert u_{j}^{\text{in}}\right\rVert}_{L^{2}}$ holds for all $t\in\mathbb{R}$. The result follows from an application of the reverse triangle inequality as
@@
Proposition˜3.5 can be adjusted without change to allow $u_{1}$ and $u_{2}$ to be initialized with periods $\frac{L}{k_{1}}$ and $\frac{L}{k_{2}}$, respectively. The only changes in this case are that $u_{j}$ remain $\frac{L}{k_{j}}$ periodic for all $t\in\mathbb{R}$ in (i) and $M<\min\{\mathinner{\!\left\lvert k_{1}\right\rvert},\mathinner{\!\left\lvert k_{2}\right\rvert}\}$ is required for (ii).
@@
The proof of this follows straightforwardly from Proposition˜3.5. Assume, by way of contradiction, that the KdV equation enjoys the finitely many determining Fourier modes property in $L^{2}$. That is, there exists some $M\in\mathbb{N}$ such that for any two solutions $u_{1}$ and $u_{2}$ to Equation˜19 that if
@@
We simply use $k=M+1$, and consider $u_{1}$ and $u_{2}$ to be any two solutions to Equation˜19 such that $u_{1}$ and $u_{2}$ both are smooth enough and have zero spatial mean, are $\frac{L}{k}$ periodic initially, and $\mathinner{\!\left\lVert u_{1}^{in}\right\rVert}_{L^{2}}\neq\mathinner{\!\left\lVert u_{2}^{in}\right\rVert}_{L^{2}}$. Two such solutions must exist as, e.g., one could take $u_{1}$ to be the zero solution and $u_{2}$ to be the solution generated from the initial data $u_{2}^{in}=\cos(\frac{2\pi}{L}kx).$
@@
The proof of the above can be adapted almost without change to other non-dissipative equations with a similar nonlinearity and conservation of the $L^{2}$ norm or a similar quantity. We will show in Section˜4 that a similar statement works for the 2D incompressible Euler equations.
@@ 3.2. Computational Results
In this section we will detail the numerical methods we utilize to simulate Equations˜19 and 20. For this computational study we implemented the KdV equation on $\mathbb{R}$, subject to periodic B.C. with period 2 using pseudo-spectral methods. Specifically, we implement the equations using an integrating factor method to compute the linear term exactly with a fourth-order Runge-Kutta scheme for the timestepping with the nonlinear term computed explicitly. For an overview of integrating factor schemes see e.g. [55, 78] and the references contained within. Our particular choice of numerical scheme, which is given Algorithm˜1, was chosen as it allows for larger timesteps than traditional schemes for KdV, such as the Zabusky-Kruskal explicit leapfrog scheme seen in the literature [80].
In the simulations of Equation˜19, depicted below, we used different spatial resolutions depending on the particular choice of initial profile and dispersion coefficient, $\delta^{2}$. We ensured in our simulations that $\Delta t\lesssim{\Delta x}^{2}$, which was consistent with the implementation of this timestepping scheme in chapter 10 of [78] and appeared to be numerically stable for all spatial resolutions we used in this work. Here $\Delta t=t_{n+1}-t_{n}$ and $\Delta x=\frac{2}{N}$ represent the temporal and spatial resolution scales with $N$ indicating the spatial resolution, given in powers of $2$. While we do not explicitly state the spatial resolution for each simulation, one can see in the spectral plots, i.e. in Figures˜4, 5, 6, 7, 8 and 9, that the x-axis covers $[1,N/2]$. NN ranges between $2^{7}$ and $2^{11}$, depending on the amount of Fourier modes required to accurately simulate a given solution. $N$ is chosen specifically to ensure that the solution is well resolved, which here means that Fourier coefficients of the reference solution decay to machine precision (given by $\varepsilon=2.216\times 10^{-16}$) before the dealiasing cutoff, i.e. the vertical red line featured in Figures˜4, 5, 6, 7, 8 and 9. We utilize 2/3 truncation for dealiasing to accurately calculate the nonlinear term, i.e., we set the last 1/3 of the Fourier modes to zero to prevent aliasing effects. Additionally, while the red vertical line on spectral plots (see Figures˜9(b), 9(d), 7(b) and 8(b) denotes the dealiasing cutoff, the vertical blue line denotes the observational window cutoff, that is, the Fourier modes to the left of and including the blue vertical line are observed and used in the computation of the nudging term.
@@
The nudging term is implemented numerically in Algorithm˜1 as an explicit term. This implementation of the feedback-control term that we used induces a CFL constraint required for the numerical scheme to remain stable, $\mu\lesssim\frac{2}{\Delta t}$. This CFL constraint can be removed by using an implicit formulation of the nudging term, which we do not consider in this work. It is worth noting that the nudging term cannot be implemented numerically using multi-stage methods, such as RK4, due to the need to interpolate in time (see e.g. [67]). The simulations featured in this work primarily use a value of $\mu=100$, which was found to be sufficient for this study. For an in-depth examination of the role the value of the nudging parameter plays in convergence, see [15].
To test the effectiveness of nudging on the KdV equation Equation˜19 we used an “identical twin” experimental design. This experimental design is standard in the validation of data assimilation methods, where we first initialize a reference solution. This solution will be treated as the true solution which we will attempt to recover using observations. The observations will be given by the lowest $M$ Fourier modes. The observations are considered to be perfect with no distortions (i.e., no errors in the measurements) or stochastic noise.
It is worth mentioning that typically in the data assimilation literature for dissipative evolutionary equations, the equations are evolved forward in time long enough such that the solution remains in an absorbing set in a relevant phase space. That is, we run solutions long enough forward in time such that they are uniformly bounded with respect to time in $L^{2}$, $H^{1}$, or another relevant norm for all future times, see e.g., [74, 76] for a full definition of absorbing sets as well as their properties. In this study we do not evolve the equations forward in time to large times, as solutions are not dissipative and so there is not necessarily an absorbing set that attracts all bounded solutions after a long enough time. In fact, as the $L^{2}$ norm is a conserved quantity for Equation˜19, such an absorbing set cannot exist in $L^{2}$. We opted to simply evolve forward the initial profile a given amount of time, typically $t=1$, in order to ensure that the reference solution has developed interesting dynamics prior to the observational period outside the observation domain. While solutions can develop small length scales not captured by the observed modes, they typically have a finite number of active Fourier modes, with the total number of modes active at a given time dependent on the initial condition, the current time, the total momentum in the system, and the dispersion parameter, $\delta^{2}$. An active Fourier mode in this sense means that $\mathinner{\!\left\lvert\hat{u}_{k}\right\rvert}\gg\varepsilon$ or $\mathinner{\!\left\lvert\hat{u}_{k}\right\rvert}>10^{j}\varepsilon$ for some reasonable $j$, where $k\in\mathbb{Z}$ is the wavemode index and $\varepsilon$ indicates machine precision.
@@ 3.2.1. Classical KdV equation
In this section we will examine how the data assimilation nudging algorithm fails for Equation˜19. The extent to which it works appears to be determined by whether the active Fourier modes of the initial profile are contained in the annulus of the observed Fourier modes, as we will see later on in Figure˜8. We will start by considering examples such as in Proposition˜3.5 to illustrate how nudging fails.
@@
Using initial data of the form Equation˜29 we can now construct solutions to Equation˜19 which the nudging algorithm Equation˜20 explicitly fails to recover. Given a number of observed modes $M$, we can set $k_{0}\in\mathbb{Z}$ such that $\mathinner{\!\left\lvert k_{0}\right\rvert}\geq M+1$. Again, we emphasize that solutions initialized in this way (see Equation˜29) are initially $\frac{2}{k_{0}}$ periodic and remain $\frac{2}{k_{0}}$ periodic due to Proposition˜3.3. Noting that $M<k_{0}$, it follows immediately from Proposition˜3.5 that $P_{M}u(t)\equiv 0$ where $u$ is the solution of Equation˜19 generated from the initial data prescribed by Equation˜29. This shows that a fixed number of observed Fourier modes are insufficient to distinguish the zero solution from solutions initialized with suitable high frequency oscillations, and so one should not expect data assimilation to succeed. Indeed, if we choose $v$ to be the solution of Equation˜20 initialized by the identically zero solution, then it is easy to deduce that $P_{M}(u-v)\equiv 0$ holds for all time and so the solution to Equation˜20 simply corresponds to a solution to Equation˜19, albeit with a different initial condition than the reference solution.
@@
One can see computationally in Figure˜4 that nudging fails to recover the solution $u(x,t)$, of Equation˜19 generated from the initial data prescribed by Equation˜29. In Figure˜4 the solution, $v(x,t)$, of Equation˜20 is initialized with the identically zero solution. Exactly as one expects, we find that $v\equiv 0$ holds for the entire simulation. Similarly, we see in Figure˜5 the roles of $u$ and $v$ are reversed. That is, we instead initialize the solution to $v$ with Equation˜29 and the reference solution, $u$, is initialized with the identically zero solution. We see in Figure˜5 similar behavior occurs, as the feedback-control term remains zero for the entire simulation and both equations evolve as solutions of Equation˜19 with different initial conditions. While this is expected, it is nevertheless an interesting result as it shows that the initialization of the assimilated solution to Equation˜20 matters. This fact is fundamentally different to many analytical results regarding the application of the nudging algorithm to dissipative equations, where the effect of the initial condition fades given enough time, contrary to what occurs in this case.
@@
The previous counterexample illustrates that data assimilation methods cannot uniquely determine the observed reference solution when the observations are identically zero, but one may hope that more physically realistic (i.e. not identically zero) observations will yield better results. This is true to some extent, however the level of success is highly dependent on the initial profile of the reference solution and one can similarly construct solutions that exhibit high frequency oscillations that nudging cannot recover. If given an analytic solution $u$ in $L^{2}$ of Equation˜19, we can indeed construct infinitely many solutions where nudging fails to synchronize with the reference solution given a finite number of observed Fourier modes.
@@
By Parseval’s identity $\mathinner{\!\left\lVert u\right\rVert}_{L^{2}}^{2}=2\sum_{k\in\mathbb{Z}}\mathinner{\!\left\lvert\hat{u}_{k}\right\rvert}^{2}=K^{2}$. Here $\varepsilon$ is a specified constant, for computations we will use $\varepsilon\sim 10^{-16}$ to denote machine precision. The purpose of finding such an $M$ is to separate the wavemodes into active low modes and inactive high modes. At any fixed time $t>0$ such an $M$ is guaranteed to exist as $u$ is analytic and therefore the Fourier coefficients must decay exponentially fast in wavenumber. In our computations we run only for a finite amount of time, so we take $M$ for $u$ to be the supremum over all such $M$’s taken across all simulated times. Without loss of generality we will assume that observation window consists of all $M$ modes, and so essentially the entire solution $u$ is known down to machine precision. We note here that $M$ is vital to the construction we outline below, thus the examples we construct work only for finite times. This is due to the fact that we require a uniform bound on the number of active wavemodes to extend this arbitrarily far in time.
Now, we can take the given solution $u$ and use it to generate artificial solutions for which nudging appears to fail. Consider the solution generated by the following initial data:
@@
Here $c_{1}$ and $c_{2}$ are positive constants chosen to enforce that $\mathinner{\!\left\lVert\tilde{u}^{in}\right\rVert}_{L^{2}}=K$ and $c_{2}>\varepsilon$. That is, we are modifying the initial profile to take some of the momentum out of the active modes and shifting it into to wavemode $2M+k$, where $k$ is an arbitrary non-negative integer. As $k$ is arbitrary, we can construct infinitely many such solutions, all of which have the same momentum as the given solution, $u$, but with a new active frequency in the initial profile. Note that the choice of using ${u}^{in}$ is arbitrary, and one can do this exact construction using the solution of $u$ evaluated at any given time.
We performed computational tests using the reference solution, $\tilde{u}$, of Equation˜19 with initial conditions prescribed by Equation˜31. Here we took a solution, $u$, of Equation˜19 which was initialized simply with $u^{in}=\cos(\pi x)$. We found $M=50$ was suitable for the construction in Equation˜31. Once $M=50$ was determined, we simply used $\tilde{u}^{in}=\cos(\pi\cdot x)+\cos(100\cdot\pi\cdot x)$, which was renormalized to ensure that $\mathinner{\!\left\lVert\tilde{u}^{in}\right\rVert}_{L^{2}}=\mathinner{\!\left\lVert u^{in}\right\rVert}_{L^{2}}$. We see in Figure˜6 the result from applying nudging to attempt to recover the constructed reference solution $\tilde{u}$. Here the nudging solution, $v$, of Equation˜20 is initialized to be identically zero. We see in Figure˜6 that the low modes of $v$ synchronize with $\tilde{u}$ to a fixed level of precision, but the reference solution contains high frequency oscillations that $v$ never develops. The error in both the high and low modes appears approximately constant after a very brief initial decay in the observed modes. In additional tests (not depicted here) the same behavior in the $L^{2}$ error was observed up to time $t=1000$ without any significant deviations in the error. One can clearly see activity in the frequencies immediately surrounding wavemodes $k=100$ and $k=200$ that is simply not present in the nudged solution. Not only is this high frequency activity never generated, but nudging features no mechanism to generate targeted high mode oscillations. That is, if one knew a priori that the solution featured oscillations at wavemode $k=100$, but one still only had observations of the first $50$ wavemodes, nudging does not have any mechanism to encourage oscillations at any unobserved frequencies.